OpenStudy (yuii):

The equation of a line is −6x − 2y = −18? What is the x-intercept of the line? x =

OpenStudy (isaiahhlindsayy):

Slope = -6.000/2.000 = -3.000 x-intercept = 9/3 = 3 y-intercept = 9/1 = 9.00000

OpenStudy (isaiahhlindsayy):

medal?

OpenStudy (yuii):

So what's the answer again?

OpenStudy (yuii):

9.0?

OpenStudy (isaiahhlindsayy):

9/3 or 3 3 is the simplified version

OpenStudy (yuii):

Alright, thank you.